OpenStudy (anonymous):

Simplify the expression -sqrt{3}-2sqrt{25}+7sqrt{21}

OpenStudy (ajprincess):

R u quite sure that u hav written the question right?

OpenStudy (anonymous):

\[-\sqrt{3}-2\sqrt{25}+7\sqrt{21}\]

OpenStudy (ajprincess):

\(-\sqrt3-2\sqrt{25}+7\sqrt{21}\) \(=-\sqrt3-2\sqrt{5*5}+7\sqrt{21}\) \(=-\sqrt3-2*5+7\sqrt{21}\) \(=-\sqrt3-10+7\sqrt{21}\) I dnt thnk u can simplify further.